# Is uniform expressivity too restrictive? Towards efficient expressivity of GNNs

### v1

@@ Appendix D Numerical experiments
Computing and approximating queries. For $Q_{1}$, the ReLU GNN has sum-aggregation, $4$ iterations (depth of $Q$). The input and output dimensions are set for convenience to $4$ to have a constant dimension for all iterations. In particular, all features vectors of the input vertices are set to $e_{1}$, the first canonical vector of $\mathbb{R}^{4}$. The weights and biases are set as follows .
$A=\begin{pmatrix}0&0&0&0\\ -1&0&0&0\\ 0&0&0&0\\ 0&0&-1&0\end{pmatrix},B=\begin{pmatrix}1&0&0&0\\ 0&0&0&0\\ 0&1&0&0\\ 0&0&0&0\end{pmatrix},c=\begin{pmatrix}-1\\ 1\\ 0\\ 1\end{pmatrix}$
$\text{For every $t\in[4]$, and every $x,y\in\mathbb{R}^{4}$ }\quad\mathsf{comb}_{t}(x,y):=\operatorname{ReLU}(Ax+By+c)$
For $Q_{2}$, the ReLU GNN has $7$ iterations. The weights and biases are set as:
$A=\begin{pmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&1&1&0&0\\ 0&0&0&0&0&0\\ 1&0&0&0&1&0\end{pmatrix},B=\begin{pmatrix}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&0&0\end{pmatrix},c=\begin{pmatrix}0\\ 0\\ 0\\ 0\\ -1\\ 0\\ -1\end{pmatrix}$

### v2

@@ Appendix D Numerical experiments
Computing and approximating queries. For $Q_{1}$, the ReLU GNN has $4$ iterations with $A$, $B$ and $c$ given by
$A=\begin{pmatrix}0&0&0&0\\ -1&0&0&0\\ 0&0&0&0\\ 0&0&-1&0\end{pmatrix},B=\begin{pmatrix}1&0&0&0\\ 0&0&0&0\\ 0&1&0&0\\ 0&0&0&0\end{pmatrix},C=\begin{pmatrix}-1\\ 1\\ 0\\ 1\end{pmatrix}.$
Note that since we are in the unicoloring setting, all vectors are initialized to $e_{1}$.
For $Q_{2}$, the ReLU GNN has $7$ iterations and $A$, $B$ and $c$ are given as:
$A=\begin{pmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&1&1&0&0\\ 0&0&0&0&0&0\\ 1&0&0&0&1&0\end{pmatrix},B=\begin{pmatrix}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&0&0\end{pmatrix},C=\begin{pmatrix}0\\ 0\\ 0\\ 0\\ -1\\ 0\\ -1\end{pmatrix}.$
Note that vertices are initialized to $e_{1}$ or $e_{2}$ depending on whether they are colored red or blue.